# Faraday Depolarization Study of a Radio Galaxy Using LOFAR Two-metre Sky Survey: Data Release 2

### v2

@@ Abstract
We present a detailed depolarization analysis of the radio galaxy ILTJ012215.21+254334.8 using polarimetric data from the LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) catalogue. This source, with RM $\sim$ 47 rad m-2 and projected linear size $\sim$ 335 kpc at $z\approx 0.05$, serves as a pilot for systematic QU-fitting of unresolved LoTSS sources, building on prior single-target studies that established the need for multi-component Faraday models in complex magneto-ionic media.
@@ 1 Introduction
When synchrotron radiation propagates through a magneto-ionic medium, or Faraday screen, the polarization angle experiences a rotation known as Faraday rotation (FR). Simultaneously, the polarized flux density may be reduced through processes collectively referred to as depolarization. Depolarization is generally divided into two main categories: external depolarization and internal depolarization (Pasetto et al., 2018). Modelling these effects requires considering the structure of the magnetic field, which may range from fully ordered to a mixture of ordered and turbulent components (Burn, 1966; Sokoloff et al., 1998; Rossetti et al., 2008; Yushkov et al., 2024).
@@
The study of Faraday depolarization in radio galaxies provides a critical window into the magneto-ionic environments both within and surrounding these powerful extragalactic sources (Strom, 1973). The goal of this study is to analyze LOFAR HBA data for a representative source from the LoTSS-DR2 catalog to test how well standard depolarization models can be constrained from unresolved fractional Stokes spectra alone, and to characterize the multi-component Faraday structure implied by the LoTSS-only data. This single-source analysis is designed as a proof of concept for a future systematic qu-fitting study of the FR-I subsample in LoTSS-DR2, where the same methodology can be applied uniformly to a statistically meaningful population. The paper is organized as follows: Section 2 introduces the key observables of radio polarization and Faraday rotation; Section 3 describes the depolarization models applied in our analysis; Section 4 outlines the main properties of the selected LoTSS DR2 source; Section 5 details the methodology; Section 6 presents the results; and Section 7 discusses the implications and provides prospects for future work.
@@ 2 Polarization and Faraday Rotation
The study of the polarization states of radio emissions enables us to analyse two important effects: Faraday rotation and Faraday depolarization (Anderson et al., 2015; Piras, 2024). Faraday rotation is caused by the rotation of the plane of polarization of electromagnetic waves as they journey across vast cosmic distances interacting with magnetized medium. Crucially, the extent of this rotation directly encodes vital information about the magneto-ionic medium—the ionized gas and embedded magnetic fields. This makes Faraday rotation an exceptionally powerful tool for probing magnetic fields not just within galaxies, but also in the vast and tenuous intergalactic medium, providing unique insights into the large-scale magnetization of the Universe that cannot be obtained through other means (Brentjens and de Bruyn, 2005). To accurately characterize the Faraday rotation for these distinct regions of polarized emission, the concept of Faraday depth $(\varphi)$ is utilized. Faraday depth $\varphi$ is defined as the integral of thermal electron density $n_{e}$ (cm-3) and parallel magnetic field $B_{\parallel}$ (µG) from the observer to a point along the line of sight, with $ds$ positive towards the source (Ferrière et al., 2021):
$\varphi=0.81\int_{0}^{s}n_{e}B_{\parallel}\,\mathrm{d}s\quad[\mathrm{rad\,\,m^{-2}}]$ (3)
@@ 3 Depolarization
(Burn, 1966), so that even intrinsically 100% polarized emission is exponentially suppressed once $2\sigma_{\text{RM}}^{2}\lambda^{4}\gg 1$. For the LoTSS HBA band (120–168 MHz, $\lambda^{2}\approx 3.2--6.3\,\text{m}^{2}$), this implies that components with $\sigma_{\text{RM}}\gtrsim 5-10\,\text{rad m}^{-2}$ are already heavily depolarized across most of the band, and for $\sigma_{\text{RM}}\gtrsim 20\,\text{rad m}^{-2}$ they would be effectively undetectable at LoTSS sensitivity even if intrinsically strongly polarized.
@@ 4 The LoTSS DR2 Catalogue and the Selected Source
In this pilot study, we therefore focus on ILTJ012215.21-254334.8, which satisfies these criteria. It has an estimated luminosity of $1.65\times 10^{24}$ W Hz-1 at 144 MHz, RM of $47.15\pm 0.05$ rad m-2, and a projected linear size of 335 kpc, i.e., very close to the LoTSS-DR2 median linear size of $\sim$ 300 kpc (O’Sullivan et al., 2023). he source has an angular size of $\sim 25^{\prime\prime}$, comparable to the 20” LoTSS beam, so the extracted spectrum averages over the entire source. With an estimated luminosity of $1.65\times 10^{24}\,\mathrm{W\,Hz^{-1}}$ at 144 MHz (below the classical FR break of $10^{25}\,\mathrm{W\,Hz^{-1}}$), this is classified as an FR-I radio galaxy (Fanaroff and Riley, 1974). The polarization fraction of approximately $3\%$ and RM are typical of the catalogued sources, so it can be regarded as representative of the dominant population of low-luminosity FR-I-like radio galaxies in LoTSS-DR2. By analysing this object in detail, the intention is to (a) validate the qu-fitting setup, priors, and instrumental-leakage treatment under realistic LoTSS conditions, and (b) identify which depolarization parameters and model classes are actually constrained by the data for a typical source. These lessons will directly inform the design and interpretation of a forthcoming study of the full DR2 FR-I sample.
@@
Here, RM is the Rotation Measure; P is the linear polarization; $p_{0}$ is the fractional polarization; I is the total intensity; $L_{144}$ is the luminosity at 144 MHz; $\ell$ is the projected linear size; and z is the redshift. The $\sim$ 335 kpc projected size implies a substantial path length through the host and cluster magneto-ionic medium, enhancing sensitivity to line-of-sight Faraday gradients and making multi-component depolarization more likely than for compact sources.
@@ 5.1 Depolarization models
1ED 1 External Faraday Dispersion 9
1T+1ED Faraday Thin + 1 External Faraday Dispersion 8 + 9
1T+2ED Faraday Thin + 2 External Faraday Dispersion 8 + 9 + 9
1Tk+1ED Faraday Thick + 1 External Faraday Dispersion 10 + 9
1Tk+2ED Faraday Thick + 2 External Faraday Dispersion 10 + 9 + 9
@@ 5.2 The priors
Model comparison follows the Kass and Raftery (1995) scale for $2\Delta\ln Z$, where values <2<2 are barely worth mentioning, values between $2$ and $6$ indicate positive evidence, values between $6$ and $10$ indicate strong evidence, and values >10>10 indicate very strong evidence against the alternative. Thus, $\Delta\ln Z>10$ decisively favours the reference model. Here, $2\Delta\ln Z>10$ indicates very strong evidence for $1T+2ED$ over alternatives (Kass and Raftery, 1995)).
@@ 6 Results
Figure 4 presents the de-rotated polarization angle $\tilde{\psi}(\lambda^{2})=\psi(\lambda^{2})+47.15\,\lambda^{2}$, where the LoTSS-DR2 catalogue RMcat = -47.15 rad m-2 has been used for de-rotation. A single thin Faraday screen at this RM would produce a flat $\tilde{\psi}=0^{\circ}$ line. Instead, the data exhibit significant scatter and curvature around this reference.
@@ Acknowledgments
SSP would like to thank Banwarilal Bhalotia College, Asansol, under Kazi Nazrul University, for facilitating the M.Sc. Physics program. This research forms part of the M.Sc. thesis work conducted in the final semester. The authors appreciate ChatGPT for helping with AI-driven copy editing and improving the manuscript’s clarity and style. AG acknowledges the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board (SERB), a statutory body under the Department of Science and Technology (DST), Government of India. AG would also like to thank IUCAA, Pune, for their support through the Associateship Programme and for providing access to their computational facilities. We would like to sincerely thank the reviewers for their thoughtful and constructive feedback on our manuscript.

### v3

@@ Abstract
We present a detailed depolarization analysis of the radio galaxy ILTJ012215.21+254334.8 using polarimetric data from the LOFAR Two-metre Sky Survey (LoTSS) Data Release 2 (DR2) catalogue. This source, with RM $\sim$ 47 rad m-2 and projected linear size $\sim$ 335 kpc at $z\approx 0.05$, serves as a pilot for systematic qu-fitting of mostly unresolved LoTSS sources, building on prior single-target studies that established the need for multi-component Faraday models in complex magneto-ionic media.
@@ 1 Introduction
When synchrotron radiation propagates through a magneto-ionic medium, or Faraday screen, the polarization angle experiences a rotation known as Faraday rotation (FR). Simultaneously, the polarized flux density may be reduced through processes collectively referred to as depolarization. Depolarization is generally divided into two main categories: external depolarization and internal depolarization (Pasetto et al., 2018). Modelling these effects requires considering the structure of the magnetic field, which may range from fully ordered to a mixture of ordered and turbulent components (Burn, 1966; Sokoloff et al., 1998; Rossetti et al., 2008; Yushkov et al., 2024). The study of Faraday depolarization in radio galaxies provides a critical window into the magneto-ionic environments both within and surrounding these powerful extragalactic sources (Strom, 1973).
@@
The paper is organized as follows: Section 2 introduces the key observables of radio polarization and Faraday rotation; Section 3 describes the depolarization models applied in our analysis; Section 4 outlines the main properties of the selected LoTSS DR2 source; Section 5 details the methodology; Section 6 presents the results; and Section 7 discusses the implications and provides prospects for future work.
@@ 2 Polarization and Faraday Rotation
The study of the polarization states of radio emissions enables us to analyse two important effects: Faraday rotation and Faraday depolarization (Anderson et al., 2015; Piras, 2024). Faraday rotation is caused by the rotation of the plane of polarization of electromagnetic waves as they journey across vast cosmic distances interacting with magnetized medium. Crucially, the extent of this rotation directly encodes vital information about the magneto-ionic medium—the ionized gas and embedded magnetic fields. This makes Faraday rotation an exceptionally powerful tool for probing magnetic fields not just within galaxies, but also in the vast and tenuous intergalactic medium, providing unique insights into the large-scale magnetization of the Universe that cannot be obtained through other means (Brentjens and de Bruyn, 2005). To accurately characterize the Faraday rotation for these distinct regions of polarized emission, the concept of Faraday depth $(\varphi)$ is utilized. Faraday depth is defined as
$\varphi(s)=0.81\int_{0}^{s}n_{e}(s^{\prime})\,B_{\parallel}(s^{\prime})\,ds^{\prime}\,\,\left[\text{rad}\,\text{m}^{-2}\right],$
where $s$ is the physical path length coordinate measured from the observer ($s=0$) towards the source, $n_{e}$ is the thermal electron density in $\text{cm}^{-3}$, and $B_{\parallel}$ is the component of the magnetic field in $\mu\text{G}$ parallel to the line of sight, taken to be positive when pointing towards the observer (Ferrière et al., 2021).
@@ 3 Depolarization
(Burn, 1966), so that even intrinsically 100% polarized emission is exponentially suppressed once $2\sigma_{\text{RM}}^{2}\lambda^{4}\gg 1$. For the LoTSS HBA band (120–168 MHz, $\lambda^{2}\approx 3.2\,\rm{m}^{2}\,-\,6.3\,\rm{m}^{2}$), this means that components with $\sigma_{\rm RM}\gtrsim 0.1\,-\,0.2\,\,\text{rad}\,\text{m}^{-2}$ are already heavily depolarized ($|p|/p_{0}\lesssim 10^{-2}$) across most of the band, and for $\sigma_{\rm RM}\gtrsim 0.3\,-\,0.5\,\,\text{rad}\,\text{m}^{-2}$ they would be effectively undetectable at LoTSS sensitivity ($|p|/p_{0}\ll 10^{-3}$) even if intrinsically strongly polarized.
@@ 4 The LoTSS DR2 Catalogue and the Selected Source
[…]
@@
Here, RM is the Rotation Measure; P is the linear polarization; $p_{0}$ is the fractional polarization; I is the total intensity; $L_{144}$ is the luminosity at 144 MHz; $\ell$ is the projected linear size; and z is the redshift. The total projected linear size of $\sim 335\,\text{kpc}$ refers to the full jet–lobe structure seen in Stokes I. In this work, however, we model only the polarization from the single $\arcsec20\arcsec$ beam at the polarized peak, i.e., from the core/inner-jet region, and not from the extended, low-surface-brightness jets.
@@ 5.1 Depolarization models
1ED 1 External Faraday Dispersion 8
1T+1ED Faraday Thin + 1 External Faraday Dispersion 7 + 8
1T+2ED Faraday Thin + 2 External Faraday Dispersion 7 + 8 + 8
1Tk+1ED Faraday Thick + 1 External Faraday Dispersion 9 + 8
1Tk+2ED Faraday Thick + 2 External Faraday Dispersion 9 + 8 + 8
@@ 5.2 The priors
Model comparison follows the Kass and Raftery (1995) scale for $2\Delta\ln Z$, where values <2<2 are barely worth mentioning, values between $2$ and $6$ indicate positive evidence, values between $6$ and $10$ indicate strong evidence, and values >10>10 indicate very strong evidence against the alternative. Thus, $2\Delta\ln Z>10$ decisively favours the reference model. Here, $2\Delta\ln Z>10$ indicates very strong evidence for $1T+2ED$ over alternatives (Kass and Raftery, 1995)).
@@ 6 Results
Figure 4 presents the de-rotated polarization angle $\tilde{\psi}(\lambda^{2})=\psi(\lambda^{2})+47.15\,\lambda^{2}$, where the LoTSS-DR2 catalogue RMcat = -47.15 rad m-2 has been used for de-rotation. Here, the raw polarization angles were first unwrapped and then de-rotated with $\pm$180 unwrapping. A single thin Faraday screen at this RM would produce a flat $\tilde{\psi}=0^{\circ}$ line. Instead, the data exhibit significant scatter and curvature around this reference.
@@ Acknowledgments
SSP would like to thank Banwarilal Bhalotia College, Asansol, under Kazi Nazrul University, for facilitating the M.Sc. Physics program. This research forms part of the M.Sc. thesis work conducted in the final semester. The authors appreciate ChatGPT for helping with AI-driven copy editing and improving the manuscript’s clarity. AG acknowledges the financial support received through the SERB-SURE grant (SUR/2022/000595) from the Science and Engineering Research Board (SERB), a statutory body under the Department of Science and Technology (DST), Government of India. AG would also like to thank IUCAA, Pune, for their support through the Associateship Programme and for providing access to their computational facilities. We would like to sincerely thank the reviewers for their thoughtful and constructive feedback on our manuscript.